OpenStudy (anonymous):

4m + 10 = 14 − 2m I tried to figure this out but i can't can some one please explain to me?

OpenStudy (studygurl14):

Your goal is to get m alone on what side of the equation. To do that, you have to rearrange terms in the equation. You can move terms from one side of the equal sign to the other by doing the opposite of that term.

OpenStudy (anonymous):

Are you trying to factor it ?

OpenStudy (anonymous):

When i get it alone it doesn't equal what the 4 answers are.

OpenStudy (anonymous):

The answers are 4 6 3/2 2/3

OpenStudy (studygurl14):

4m + 10 = 14 - 2m You can move the -2m to the other side by adding 2m because (-2m + 2m = 0) 4m + 10 = 14 - 2m 4m + 2m + 10 = 14 + 0

OpenStudy (studygurl14):

4m and 2m can be combined into 6m 6m + 10 = 14

OpenStudy (studygurl14):

next, move the 10 over to the right side of the equation. Do you know how to do that?

OpenStudy (anonymous):

yes i do now its a two step equation

OpenStudy (anonymous):

But i still can't' get the answer because i tried that already.

OpenStudy (studygurl14):

Let's take this step by step. What's the next step?

OpenStudy (anonymous):

do the opposite of addition correct? so 14-10

OpenStudy (studygurl14):

correct. And you would get?

OpenStudy (anonymous):

4

OpenStudy (studygurl14):

right. So you are left with what equation?

OpenStudy (anonymous):

6m = 4

OpenStudy (studygurl14):

correct. now what? remember, you want to isolate the m. right now, the m is being multiplied by 6. What is the opposite of multiplying by 6?

OpenStudy (anonymous):

dividing

OpenStudy (studygurl14):

correct. So what would be the result?

OpenStudy (anonymous):

Which then it goes into a repeating .6

OpenStudy (anonymous):

I have trouble converting into fractions from decimals.

OpenStudy (anonymous):

So i might mess the other part up.

OpenStudy (studygurl14):

You don't need to convert from fraction to decimal.

OpenStudy (anonymous):

4 6 3/2 2/3

OpenStudy (anonymous):

Are the answer choices

OpenStudy (studygurl14):

Okay, so you have m = 4/6 right?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

you can simplify that fraction without turning it into a decimal. Do you know how?

OpenStudy (anonymous):

Yes

OpenStudy (studygurl14):

okay, so what do you get?

OpenStudy (anonymous):

2/3

OpenStudy (studygurl14):

Right! That's your answer

OpenStudy (anonymous):

Thank you i am also confused with one more.

OpenStudy (studygurl14):

You're welcome. Feel free to post your other question after closing this one and link me to it. :)

OpenStudy (studygurl14):

Do you know how to give medals to people who help you?

OpenStudy (anonymous):

yes i do

OpenStudy (studygurl14):

:)